### v1

# A Survey of Quantized Graph Representation Learning: Connecting Graph Structures with Large Language Models

Abstract
Recent years have witnessed rapid advances in graph representation learning, with the continuous embedding approach emerging as the dominant paradigm. However, such methods encounter issues regarding parameter efficiency, interpretability, and robustness. Thus, Quantized Graph Representation (QGR) learning has recently gained increasing interest, which represents the graph structure with discrete codes instead of conventional continuous embeddings. Given its analogous representation form to natural language, QGR also possesses the capability to seamlessly integrate graph structures with large language models (LLMs). As this emerging paradigm is still in its infancy yet holds significant promise, we undertake this thorough survey to promote its rapid future prosperity. We first present the background of the general quantization methods and their merits. Moreover, we provide an in-depth demonstration of current QGR studies from the perspectives of quantized strategies, training objectives, distinctive designs, knowledge graph quantization, and applications. We further explore the strategies for code dependence learning and integration with LLMs. At last, we give discussions and conclude future directions, aiming to provide a comprehensive picture of QGR and inspire future research.
1 Introduction
Graph representation usually involves transforming nodes, edges, or structures within graph data to low-dimensional dense embeddings, which is called continuous graph representation as shown in Figure 1 (a). The learned representation is obligated to preserve both node attributes and topological structure Zhu et al. (2020). This field has seen substantial development and expansion in recent years. The emergence of numerous advanced technologies has invariably drawn significant attention within the graph community, becoming focal points of extensive studies, e.g., the well-known node2vec Grover and Leskovec (2016) model, graph neural networks (GNNs) Kipf and Welling (2017), and self-supervised graph learning Liu et al. (2022). They have achieved great empirical success in graph-related domains, including bioinformatics, social networks, recommendation systems, and anomaly detection Xia et al. (2021); Liu et al. (2023c).
Although continuous graph representation successfully extracts rich semantic information from the input graph to adapt to a wide range of machine learning and deep learning tasks, concerns may arise regarding the embedding efficiency, interpretability, and robustness Wang et al. (2024). Specifically, with the great growth of graph scales, there will be a linear increase in embedding parameters, becoming a considerable challenge in graph learning contexts involving millions or even more nodes. Besides, continuous embeddings are generated in a black-box fashion and the meaning of the overall representation or any one dimension is unknowable, leading to a lack of interpretability. Moreover, these continuous representations are typically task-specific and show no significant abstraction at the representation level, making them not robust enough to generalize to a variety of practical tasks as effectively as LLMs.
Interestingly, in real life, discrete features are often preferred for representation. For example, as shown in Figure 1 (b), descriptions affiliated to an individual are often a combination of discrete and meaningful features, such as age (e.g., 18), gender (male or female), and place of residence (e.g., Beijing). This scheme is simple and easy to understand. Based on this intuition, quantized representation came into being, which is now a hot topic in the computer vision domain that can facilitate image generation by combining with Transfomer or LLMs Esser et al. (2021); Lee et al. (2024). Inspired by general quantized representation techniques, quantized graph representation (QGR) learning has recently been proposed to address the aforementioned shortcomings of continuous graph learning. The core idea is to learn discrete codes for nodes, subgraphs, or the overall graph structures, instead of employing continuous embeddings. As shown in Figure 1 (b), a node can be represented as discrete codes of (male, 18, Beijing, Ph.D.). In QGR methods, codes can be learned by differentiable optimization with specific graph targets or self-supervised graph learning. It has the capacity to learn more high-level representations due to a compact latent space, which would enhance the representation efficiency, interpretability, and robustness. Recall that when faced with graph-text scenarios, incorporating continuous embeddings into advanced LLMs is inherently challenging and prone to introducing the loss of semantic information. In contrast, due to the consistency with the discrete nature of natural language, the output of QGR would be seamlessly integrated into LLMs, like early-fusion strategy Team (2024). Considering that this emerging paradigm is still in its initial stage but carries substantial potential in the era of LLMs, we undertake this thorough survey. It encompasses various viewpoints, including different graph types, quantized strategies, training objectives, distinctive designs, applications and so forth. We also provide discussions and conclude future directions, intending to motivate future investigation and contribute to the advancement of the graph community.
To start with, we first briefly give the background in Section 2. Then, we provide a comprehensive introduction to QGR frameworks from multiple perspectives in Section 3. Next, we present the code dependence learning as well as the integration with LLMs in Section 4 and Section 5, respectively. Finally, we give the discussions and future directions in Section 6.
2 Background
2.1 Merits of Quantized Graph Representation
Compared to continuous representation methods, QGR offers a series of distinct advantages that fertilize its applications.
Embedding Parameter Efficiency.
As large-scale graphs become more prevalent, there is a corresponding substantial increase in the representation parameters, which necessitates significant memory usage for storage and substantial computational expenses Galkin et al. (2022). Supposing as the number of nodes and as the dimensions, the required embedding parameters would be . In the quantized settings, the total count of parameters would reduce to , where and denote the length of the codeword set and the number of codes assigned to each node. For example, if there are nodes with 1024 feature dimensions, 1024 codewords, and 8 assigned codes, it would be a 113-fold reduction in required embedding parameters.
Explainability and Interpretability.
The discrete nature of QGR brings the explainability and interpretability for embeddings and reasoning, as shown in Figure1 (b). Each code would be assigned to the practical item by specific designs for direct interpretability, e.g., Dr. E Liu et al. (2024b) introduces a language vocabulary as the code set and each node in graphs can be represented as the permutation of explainable language tokens.
Robustness and Generalization.
The discrete tokens of the QGR are more robust and generalizable, which usually utilizes the self-supervised strategies to learn both local-level and high-level graph structures (e.g., edge reconstruction and DGI Velickovic et al. (2019)) as well as semantics (e.g., feature reconstruction). Moreover, the acquired tokens can readily adapt to downstream applications, either directly or indirectly Yang et al. (2024).
Seamless Integration with NLP Models.
The swift advancement in natural language processing (NLP) techniques, particularly LLMs, has sparked an increased interest in employing NLP models to resolve graph tasks. However, the inherent representation gap between the typical graph structure and natural language poses a significant challenge to their seamless and effective integration. Graph quantization, by learning discrete codes that bear resemblance to natural language forms, can facilitate this integration directly and seamlessly Lin et al. (2025a); Liu et al. (2024b).
2.2 General Quantization Methods
The mainstream quantization methods can be roughly categorized into product quantization, vector quantization, finite scalar quantization, and anchor selection and assignment.
Product quantization (PQ).
The core idea of PQ Jegou et al. (2011) is to divide a high-dimensional space into multiple low-dimensional subspaces and then quantize them independently within each subspace. Specifically, it splits a high-dimensional vector into multiple smaller subvectors and quantizes each subvector separately. In the quantization process, each subvector is mapped to a set of a finite number of center points by minimizing the quantization error.
Vector Quantization (VQ).
To effectively learn quantized representations in a differentiable manner, the VQ strategy van den Oord et al. (2017); Esser et al. (2021) is proposed and has emerged as the predominant method in the field. The core idea involves assigning the learned continuous representations to the codebook’s index and employing the Straight-Through Estimator Bengio et al. (2013) for effective optimization. The formulation is as follows. , , denote the encoder, decoder, and quantization process, respectively. is the codebook representation with dimension , which corresponds to discrete codewords . is the graph, where is the node set with size and is the edge set. The presence of or is not guaranteed, where is the input feature of a graph and is the label set of each node. is the latent representation after encoder and is the reconstruction feature after quantization and decoder, i.e., , .
Specifically, the VQ strategy yields a nearest-neighbor lookup between latent representation and prototype vectors in the codebook for quantized representations:
indicating that node is assigned codeword . Further, the model can be optimized by the Straight-Through Estimator:
where sg represents the stop-gradient operator. is a hyperparameter to balance and is usually set to 0.25. In this loss, the first codebook loss encourages the codeword vectors to align closely with the encoder’s output. The second term of the commitment loss aids in stabilizing the training process by encouraging the encoder’s output to adhere closely to the codebook vectors, thereby preventing excessive deviation.
Using VQ techniques, a continuous vector can be quantized to the counterpart of a discrete code. However, there would be distinctions and differences in the vector pre and post-quantization. To address it, Residual Vector Quantization (RVQ) Lee et al. (2022) is proposed to constantly fit residuals caused by the quantization process. The whole process can be seen in Algorithm 1. RVQ is actually a multi-stage quantization method and usually has multiple codebooks. The codebook stores the residual values of the quantized vector and the original vector at each step, and the original vector can be approximated infinitely by quantizing the residual values at multiple steps.
Finite Scalar Quantization (FSQ).
Because of the introduction of the codebook, there would be complex technique designs and codebook collapse in VQ techniques. Thus, FSQ Mentzer et al. (2024) introduces a very simple quantized strategy, i.e., directly rounding to integers rather than explicitly introducing parameters of the codebook. FSQ projects the hidden representations into a few dimensions (usually fewer than 10) and each dimension is quantized into a limited selection of predetermined values, which inherently formulates a codebook as a result of these set combinations. It is formulated as:
where is the rounding operation and is the number of unique values. The process of rounding simply adjusts a scalar to its nearest integer, thereby executing quantization in every dimension. is a sigmoid or tanh function. Its loss is calculated by also using the Straight-Through Estimator:
FSQ has the advantages of better performance, faster convergence and more stable training, particularly when dealing with large codebook sizes.
Anchor Selection and Assignment (ASA).
It is typically unsupervised and often employs prior strategies to identify informative nodes within the graph, e.g., the Personalized PageRank Page (1999) and node degree. Consequently, each node can be attributed to a combination of anchors that exhibit structural or semantic relationships, like using the shortest path. This form of quantization designates the code as a real node, unlike the aforementioned three methods.
3 Quantized Graph Learning Framework
We summarize the main studies of QGR in Table 1 from the primary perspectives of graph types, quantization methods, training strategies, and applications. Additionally, we also highlight if there are in pipeline and self-supervised manners, as well as whether they learn code dependence and integrate with language models.
The general process of QGR is illustrated in Figure 2, which comprises the encoder , decoder , and quantization process . The encoder is to model the original graph structure into latent space, where MLPs He et al. (2020) and GNNs Xia et al. (2023) are usually utilized. The decoder also commonly utilizes GNNs for structure embedding and the quantization process implements the strategies outlined in Section 2.2. In particular, Bio2Token Liu et al. (2024a) utilizes Mamba Gu and Dao (2023) as both encoder and decoder. UniMoT Zhang et al. (2024) leverages the pre-trained MoleculeSTM molecule encoder Liu et al. (2023a) to connect the molecule graph encoder and causal Q-Former Li et al. (2023b). In the following sections, we will delve into the details of the model design.
Model | one/p | Self-Sup | Graph Type | Quantization | Training | Application | Dependence | w/ LM |
[…]
3.1 Quantized Strategies
Except for SNEQ He et al. (2020) and d-SNEQ He et al. (2023), the majority of research tends to employ VQ-related strategies. Beyond the general techniques in VQ, Dr.E Liu et al. (2024b) utilizes RVQ in one specific layer, which is called intra-layer residue. For more effective in encoding the structural information of the central node, it involves preserving the multi-view of a graph. Specifically, between GCN layers, the inter-layer residue is carried out to enhance the representations:
where and denote the latent representations and quantized ones. is the number of learned codes for each layer. Using this process, the discrete codes of each node are generated in a gradual (auto-regressive) manner, following two specific orders: from previous to subsequent, and from the bottom layer to the top. This approach guarantees comprehensive multi-scale graph structure modeling and efficient code learning. iMoLD Zhuang et al. (2023) proposes another type of Redidual VQ, incorporating both the continuous and discrete representations, i.e., , as the final node representation.
GLAD Boget et al. (2024b) and Bio2Token Liu et al. (2024a) utilize the straightforward approach, FSQ, to facilitate the learning of discrete codes. For KG quantization, early methods such as NodePiece Galkin et al. (2022), EARL Chen et al. (2023), and RandomEQ Li et al. (2023a) have adopted the unsupervised learning approach ASA to achieve effective learning outcomes. For more condensed code representations, SSQR Lin et al. (2025a) employs the VQ strategy to capture both the structures and semantics of KGs.
3.2 Training Objectives
In this section, we will delve deeper into training details from the perspectives of node, edge, and graph levels.
Node Level.
Feature reconstruction usually serves as a simple self-supervised method for QGR learning, which is to compare the original node features with reconstructed features based on the learned codes. It mainly has the following two implementation types:
The first involves employing cosine similarity with scaled parameter , as implemented in Mole-BERT and VQGraph. The second is the mean square error as in Dr.E. NID Luo et al. (2024) utilizes GraphMAE Hou et al. (2022) for self-supervised learning, which can be viewed as a masked feature reconstruction. It involves selecting a subset of nodes, masking the node features, encoding by a message-passing network, and subsequently reconstructing the masked features with a decoder.
Edge Level.
It is a primary strategy to acquire the graph structures in QGR codes using a self-supervised paradigm. The direct link prediction, i.e., edge reconstruction, is a widely adopted technique to assess the presence or absence of each edge before and after the reconstruction process as:
where the mean square error (e.g., VQGraph) and the binary cross-entropy are utilized to optimize (e.g., Dr.E), respectively. is the prediction of probability for the existence of the edge based on the learned representations. is the ground label.
Beyond direct edge prediction, He et al. (2020) proposes high-level connection prediction between two nodes, i.e., structural connection, and semantic connection. The former aims to accurately maintain the shortest distances within the representations:
where is to calculate the distance of two nodes based on the learned codes. Considering that nodes sharing identical labels ought to be situated closer together within the embedding space, semantic connection is proposed:
where is the sample size. presents the constant semantic margin, which is set to 0 if the labels of two nodes are totally distinct, otherwise, it is set to a constant.
Graph Level.
For high-level representations, the graph modeling targets are proposed. For instance, graph contrastive learning is a widely utilized method for modeling relationships between graph pairs:
where is the dataset that contains all graphs and is the graph representation based on . and can be seen as positive samples for the target graph, which are usually transformed by the target graph . For example, Mole-BERT Xia et al. (2023) utilizes different masking ratios (e.g., 15% and 30%) for molecular graphs to obtain and . is contrastive samples, typically derived from the sampled batch. GraphCL You et al. (2020) is also a well-known framework of this kind based on graph augmentations, including node dropping, edge perturbation, attribute masking, and subgraph, which is used by NID Luo et al. (2024).
Furthermore, motivated by a straightforward self-supervised learning framework that identifies varied augmentation views as similar positive pairs, iMoLD Zhuang et al. (2023) considers the learned invariant and global aspects (invariant plus spurious) as positive pairs, aiming to enhance their similarity. An MLP-based predictor is introduced to transform the output of one view and align it with the other. The whole process can be interpreted as invariant learning. Deep Graph Infomax (DGI) Velickovic et al. (2019) can also be used as a high-level QGR learning method, which is a contrastive approach that compares local (node) representations with global (graph or sub-graph) ones. GQT Wang et al. (2024) utilize both high-level DGI Velickovic et al. (2019) and node-level GraphMAE2 Hou et al. (2023) as training targets.
3.3 Distinctive Designs
Codebook Design.
Beyond randomly setting the codebook for updating, there are several specifically designed strategies for domain knowledge adaption and generalization. For example, Dr.E Liu et al. (2024b) sets the vocabulary of LLaMA as the codebook, realizing token-level alignment between GNNs and LLM. In this way, each node of graphs can be quantized to a permutation of language tokens and then can be directly input to LLMs to make predictions. When modeling molecular graphs, Mole-BERT Xia et al. (2023) categorizes the codebook embeddings into various groups, each representing a distinct type of atom. For instance, the quantized codes of carbon, nitrogen and oxygen are confined to different groups.
Training Pipelines.
Aside from the single-stage learning process applied to QGR and particular tasks, some methods incorporate multiple stages into the training framework. GQT Wang et al. (2024) modulate the learned codes through hierarchical encoding and structural gating, which are subsequently fed into the Transformer network and aggregate the learned representations through an attention module. Based on the learned quantized representation, NID Luo et al. (2024) trains an MLP network for downstream tasks, such as node classification and link prediction.
3.4 KG Quantization
Recently, several KG quantization methods have been proposed for effective embedding, responding to the increasing demand for larger KGs. Unlike the techniques employed in the HOG or HEG setting, these methods tend to leverage an unsupervised paradigm for quantization. NodePiece Galkin et al. (2022), EARL Chen et al. (2023), and random entity quantization (RandomEQ for short) Li et al. (2023a) are all in such a framework. They first select a number of entities as anchors and then utilize the structural statistical strategy to match them for each entity. Specifically, NodePiece employs metrics such as Personalized PageRank Page (1999) and node degree to select certain anchor entities and subsequently view top- closest anchors as entity codes. EARL and RandomEQ sample 10% entities as anchors and then assign soft weights to each anchor according to the similarity between connected relation sets. Given the limitations of these methods in thoroughly capturing the structure and semantics of KGs, SSQR presents a self-supervised approach that employs learned discrete codes to reconstruct structures and imply semantic text, offering more holistic modeling with only 16 quantized codes.
3.5 Application Scenarios
General Graph Tasks.
Based on the learned quantized codes, many general graph tasks can be addressed, for example, node classification, link prediction, graph classification, graph regression, and graph generation. In addition, SNEQ He et al. (2020) can be employed for node recommendation that ranks all nodes based on a specific distance metric and suggests the nearest node. d-SNEQ He et al. (2023) is further utilized for path prediction that predicts the path (e.g., the shortest path) between two nodes.
Based on the learned structure-aware codes for nodes, VQGraph Yang et al. (2024) enhances the GNN-to-MLP distillation by proposing a new distillation target, namely soft code assignments, which utilizes the Kullback–Leibler divergence to make two distributions (codes distributions by GNN and MLP) be close together. This can directly transfer the structural knowledge of each node from GNN to MLP. The results show it can improve the expressiveness of existing graph representation space and facilitate structure-aware GNN-to-MLP distillation. In the KG scenarios, NodePiece Galkin et al. (2022), EARL Chen et al. (2023), RandomEQ Li et al. (2023a), SSQR Lin et al. (2025a) all can be used for KG link prediction, which is a ranking task that to predict the object entity based the given subject entity and the relation, i.e., . Using the learned codes as features, SSQR can also fix the triple classification problem, predicting the validity of the given triple .
Molecular Tasks & AI for Science.
Recently, QGR methods have been widely used for molecular tasks and AI for science scenarios, including molecular property prediction (classification and regression), molecule-text prediction, and protein & RNA reconstruction. Specifically, DGAE Boget et al. (2024a) conduct the graph generation. It first iteratively samples discrete codes and then generates the graph structures by the pre-trained decoder, which can be used in the generation for molecular graphs and has the potential for drug design, material design and protein design. LLPS Gaujac et al. (2024) shows the potential for the reconstruction of protein sequences by training a de novo generative model for protein structures using a vanilla decoder-only Transformer model. Bio2Token Liu et al. (2024a) conducts efficient representation of large 3D molecular structures with high fidelity for molecules, proteins, and RNA, holding the potential for the design of biomolecules and biomolecular complexes. UniMoT Zhang et al. (2024) unifies the molecule-text applications, including molecular classification & regression, molecule captioning, molecule-text retrieval, and caption-guided molecule generation. It demonstrates the impressive and comprehensive potentials that arise from the combination of QGR and LLMs.
4 Code Dependence Learning
Inspired by the distribution learning of the discrete codes in computer vision field Esser et al. (2021) that predicts the next code by the auto-regressive Transformer, QGR methods also implement this technique, facilitating comprehensive semantic modeling and supporting the generation tasks. Similar to BERT Devlin et al. (2019) pre-training style, MOLE-BERT adopts a strategy similar to Masked Language Modeling (MLM). It employs this method to pre-train the GNN encoder by randomly masking certain discrete codes. Subsequently, it pre-trains GNNs to predict these masked codes, a process known as Masked Code Modeling (MCM):
where is the set of masked nodes in the graph and is the quantized codes. Auto-regressive code dependence learning is another mainstream strategy (usually using Transformer) based on next token prediction:
where network is with parameter . is the length of the code sequence c. DGAE Boget et al. (2024a) splits the latent representation of each node into parts. So after the quantization, each graph can be represented as code permutation. To learn the dependence of it, the 2D Transformer Vaswani et al. (2017) is introduced to auto-regressively generate the codes, i.e., the joint probability can be .
Moreover, GLAD Boget et al. (2024b) implements diffusion bridges Liu et al. (2023b) for codes’ dependence, where Brownian motion is utilized as a non-conditional diffusion process defined by a stochastic differential equation (SDE). Based on the learned model bridge, discrete practical codes can be acquired by the iterative denoising process from the random Gaussian noise. Followed by the pre-trained decoder, these codes can be used for the graph generation. Although Dr.E does not directly or explicitly learn the dependence of the codes, it introduces intra-layer and inter-layer residuals to gradually generate the codes, which enhances the representation of sequential information as the newly generated code depends on the previously obtained codes. It can be viewed as an implicit auto-regressive manner.
5 Integration with LLMs
The quantized codes, being discrete, share a similar structure with natural language. As such, QGR methods can be seamlessly incorporated with LLMs to facilitate robust modeling and generalization as shown in Figure 3. Table 2 provides examples of tuning instructions for both Dr.E and SSQR, arranged for better understanding and intuitive interpretation.
(a) Node classification, taken from Dr.E Liu et al. (2024b). |
(b) KG triple classification, taken from SSQR Lin et al. (2025a). |
Dr.E implements token-level alignment between GNNs and LLMs. Based on the quantized codes through intra-layer and inter-layer residuals, the preservation of multiple perspectives in each convolution step is guaranteed, fostering reliable information transfer from the surrounding nodes to the central point. Moreover, by utilizing the LLMs’ vocabulary as the codebook, Dr.E can represent each node with real words from different graph hops, as shown in Table 2 (a). By fine-tuning LLaMA-2-7B Touvron et al. (2023) with the designed instruction data, the node label can be easily predicted. Different from the fixed vocabulary in Dr.E, UniMoT Zhang et al. (2024) expands the original LLMs’ vocabulary through the incorporation of learned molecular tokens from the codebook, forming a unified molecule-text vocabulary. Based on the specific instructions for molecular property prediction, molecule captioning, molecule-text retrieval, caption-guided molecule generation, LLaMA2 is fine-tuned with LoRA Hu et al. (2022), thereby bestowing upon it an impressive and comprehensive proficiency in molecule-text applications. Similarly, in KG-related scenarios, SSRQ leverages the learned discrete entity codes as additional features to directly input LLMs, helping to make accurate predictions. Both LLaMA-2-7B and LLaMA3.1-8B Dubey et al. (2024) are utilized for the KG link prediction task and triple classification task. During the fine-tuning phase, the introduction of new tokens necessitates the expansion of the LLM’s tokenizer vocabulary, which is in line with the size of the codebook, e.g., 1024 or 2048. All other elements of the network remain unaltered.
6 Discussions and Future Directions
Beyond the above methods, there are also some methods that does not use the quantized code for the final representation, such as VQ-GNN Ding et al. (2021), VQSynery Wu et al. (2024), and MSPmol Lu et al. (2024). For example, instead of employing discrete motif features as the ultimate representation, MSPmol Lu et al. (2024) utilizes the Vector Quantization (VQ) method on motif nodes within specific layers of the molecular graph. There are also benefits for the representation, which can be viewed as a paradigm of contiguous+quantized with the network, rather than the outputs. They differ from the studies in Table 1, so we do not include them in this survey.
Despite some achievements of the QGR methods, there are also some shortcomings from both methodology and application perspectives. First, the quantization process may result in the loss of some details of the original data, which could pose challenges in areas that require accurate modeling. Second, the quantization process can require complex techniques to ensure effective and efficient optimization, while pinepine’s schema and integration with LLMs add additional complexity to the overall framework. Third, compared with the wide application and success of general LLMs, the research on the integration of QGR and LLMs is still lacking. Thus, the direction of future development lies in the following four aspects.
Choice of Codeword.
For QGR learning, one can incorporate certain domain knowledge into the construction of the codebook. This knowledge might encompass semantic concepts, structural formations, and textual language. Such an approach can augment the interpretability of the whole framework, thereby facilitating a better human understanding.
Advanced Quantized Methods.
GQR can adopt the latest advanced quantized methods to enhance the efficiency and efficacy of learning graph codes, which may have already demonstrated success within the domain of computer vision. For instance, the rotation trick Fifty et al. (2024) is a proven method to mitigate the issues of codebook collapse and underutilization in VQ. SimVQ Zhu et al. (2024) simply incorporates a linear transformation into the codebook, yielding substantial improvement.
Unified Graph Foundation Model.
While the GQR methods have indeed made some notable strides, the current research primarily follows a similar approach that tackles various tasks using distinct training targets, resulting in multiple distinct models. This could restrict its applicability, particularly in the age of LLMs. Drawing inspiration from the remarkable success of unified LLMs and multimodal LLMs Zhao et al. (2023); Lin et al. (2025b), the unified graph foundation model could be realized to execute various tasks across diverse graphs through QGR, by learning unified codes for different graphs and then tuning with LLM techniques.
Graph RAG with QGR.
Retrieval-Augmented Generation (RAG) has emerged as a significant focal point for improving the capabilities of LLMs within specific fields. Graph RAG Edge et al. (2024) would be beneficial for downstream tasks in the context of LLMs, where the QGR learning would provide an effective manner to retrieve relevant nodes or subgraphs by calculating the similarity of the codes. It can be easily realized by the statistical metrics of the codes or the numerical calculations among embeddings.